Gibbs Free Energy/ Thermochemistry

1. Calculate the equilibrium constant at the temperature given

O2(g) +2F2(g) --> 2F2O (g) (T=100 C)

Delta G -79 KJ
Answer #1

T= 100.0 oC

= (100.0+273) K

= 373 K

ΔGo = -79 KJ/mol

ΔGo = -79000 J/mol

use:

ΔGo = -R*T*ln Kc

-79000 = - 8.314*373.0* ln(Kc)

ln Kc = 25.4747

Kc = 1.157*10^11

Answer: 1.16*10^11